# Watt-level ultrafast 1.75 µm laser system based on thulium-doped core and terbium-doped cladding fluoride fibers

### v1

@@ 3 Pulse compression results
Finally, we examine the tuning effect of the CFBG using a constant pumping power of 9.75 W and fix grating distance (174 mm). The FROG trace of the compressed pulse is shown in Fig. 5(a). The intensity and phase of the retrieved pulse in time and frequency domains are also shown in Fig. 5(b) and Fig. 5(c), respectively. Using an offset in the dispersion slope of 50 ps/nm2 which corresponds to GDD=7.1296 ps2 and TOD=0.0578 ps3, we measured the shortest pulse of 217 fs. The transform-limited pulse duration derived from the spectrum was 216 fs. The close match between the transform-limited and retrieved pulse duration proves that the CFBG dispersion profile was successfully compensated to that of the grating pair. Although some residual TOD can be seen in the spectral phase, we did not observe any serious phase distortion due to the nonlinear effects in the power amplifier. The compressed pulses peaked near 1.76 µm, which is blue-shifted relative to our previous system [okada2024femtosecond]. Excellent temporal quality of the compressed pulse was also achieved as the retrieved temporal profile shows high contrast. That no considerable energy is distributed on the sides of the central pulse. Therefore, we expect that multi-photon excitation now will be more efficient. The power of the pulse after the Treacy compressor was estimated to be 700 mW.
4 Summary and conclusion
@@
In summary, we have improved our previous CPA system by redesigning the amplification stage and incorporating a CFBG for tunable dispersion control in the compression stage. The output power reached 1 W. With the controllable CFBG, we were able to compress the pulse to 217 fs. The pulse was compressed with reasonable spectral quality, as the spectrogram showed independence from the pumping power, and with an improved temporal profile, as it is free from side artifacts and strong nonlinear effects. We believe this is the first all-Tm:Tb:ZBLAN amplifier system that directly generates watt-level, short-wavelength ultrafast pulses. The construction of the all-Tm:Tb:ZBLAN amplifier system offers a simpler design while still delivering high average power and pulse duration near 200 fs, making it a suitable light source for three-photon lifetime imaging microscopy experiments with red fluorescent proteins [cheng2014measurements, liu2018ex, prevedel2025three].
@@
National Institute for Physiological Sciences (24NIPS225, 25NIPS233); National Institute of Natural Sciences (OML0123008,OML012410);

### v2

@@ 3 Pulse compression results
Finally, we examine the effect of the CFBG tuning using a constant pumping power of 9.75 W and a fixed grating distance of 174 mm. We fine-tuned the CFBG by applying offsets in the dispersion slope from 90 ps/nm2 to 90 ps/nm2. The best FROG trace of the compressed pulse is shown in Figure 5(a). The intensity and phase of the pulse retrieved in the time and frequency domains are also shown in Figures 5(b) and 5(c), respectively. Applying a dispersion offset of 50 ps/nm2, equivalent to a GDD of 7.1296 ps2 and a TOD of 0.0578 ps3, resulted in a minimum pulse duration of 237 fs. The transform-limited pulse duration derived from the spectrum was 210 fs. The close match between the transform-limited and retrieved pulse duration proves that the CFBG dispersion profile was successfully compensated to that of the grating pair. An excellent temporal quality of the compressed pulse was also achieved, as the retrieved temporal profile shows high contrast. There is no considerable energy distributed on the sides of the central pulse. Therefore, we expect that multi-photon excitation now will be more efficient. Although some residual TOD can be seen in the spectral phase, we did not observe any serious phase distortion due to the nonlinear effects in the power amplifier. The retrieved spectral profile of the compressed pulses peaked near 1.76 µm, which is blue-shifted relative to our previous system [22]. It is consistent with the independently measured spectrum immediately after the Treacy compressor (recorded using a AQ6376 spectrometer), as depicted by the black solid plot overlapped in Figure 5(c). Overall, the consistency between the compressor-output spectrum and the FROG-retrieved spectrum demonstrates that the SHG-FROG accurately captures the dominant spectral content of the pulses. The pulse power after the Treacy compressor was estimated to be 700 mW.
4 Amplification and compression results for the short Tm:Tb:ZBLAN fiber
We also investigated the amplification and compression characteristics of pulses generated using a shorter gain fiber. In the new power amplifier configuration, the length of the gain medium was reduced to 1.6 m. The ion concentrations were maintained at 2500 ppm and 3000 ppm for the Tm and Tb ions, respectively, and the NA is 0.14. However, the shorter fiber had a larger core diameter of 8.5 µm compared to 7.6 µm for the 4 m fiber. All other fiber components and laser components remained similar to those with the 4 m fiber.
We anticipated that the amplified pulse spectrum would be broader when using the 1.6 m fiber compared to the 4 m fiber, because shorter fiber exhibits a broader ASE spectrum range. However, a comparison of Figure 3(c) and Figure 6(a) shows otherwise. We note that the apparent broadening in Figure 3(c) arises primarily from long-wavelength components and unsuppressed ASE during amplification. In contrast, the short fiber yields a higher signal-to-noise ratio after amplification compared to the long gain fiber. Specifically, for the 4 m fiber, the spectral component at 1.75 µm is only about 5 dB higher than that at 1.8 µm, as shown in Figure 3(c), while for the 1.6 m fiber this difference increases to nearly 20 dB, as depicted in Figure 6(a).
Similar to the 4 m fiber case, no obvious spectral narrowing or broadening was observed, indicating that nonlinear effects were negligible. The output power increased linearly with absorbed pump power up to 3.61 W, beyond which slight saturation was observed. By fitting the output power up to 3.61 W, the slope efficiency was calculated to be 15.6. At an absorbed pump power of 5.10 W, the output power reached 0.7 W and we estimated 175 nJ of single pulse energy. The slight improvement in efficiency observed with the short fiber can be attributed to its larger core diameter.
Essentially, in a small-core, long-fiber amplifier (4 m fiber), the pump light is absorbed very strongly near the input end of the fiber. This leads to a high population inversion at the beginning, but the pump is depleted quickly, causing the inversion to drop significantly as the light propagates. As a result, a large portion of the fiber toward the output can become weakly inverted or even absorbing for the signal. The extended fiber length also provides a large gain path for ASE, allowing ASE to build up and compete with the signal, which reduces overall efficiency. In contrast, a large-core, short-fiber amplifier (1.6 m fiber) operates with a lower pump intensity per unit area, but the pump absorption is spread more evenly along the fiber length. This leads to a higher average population inversion throughout the active region rather than a strong inversion localized at the input. Because the fiber is shorter, there is less opportunity for ASE growth and signal reabsorption, so more of the stored energy is efficiently transferred to the desired signal.
Finally, the amplified pulses generated by the 1.6 m fiber were compressed using the same grating-pair compressor. The GDD was reduced by varying the separation of the gratings, and the optimal distance was found to be L = 183 mm. The best FROG trace of the compressed pulse, shown in Figure 7(a), was obtained with a dispersion offset corresponding to a GDD of 7.1348 ps2 and a TOD of 0.0583 ps3. Under these conditions, the shortest measured pulse duration was 222 fs as presented in Figure 7(b). The transform-limited pulse duration derived from the spectrum was 212 fs. The close agreement between the transform-limited and retrieved pulse durations indicates that the CFBG dispersion profile was effectively compensated for by the grating pair. As depicted in Figure 7(c), we were able to compress the pulse with reasonable quality, as confirmed by a small residual spectral phase. The retrieved compressed pulses exhibit a spectral peak near 1.75 µm, in good agreement with the independently measured spectrum recorded immediately after the Treacy compressor, as shown by the solid black line spectrum overlapped in Figure 7(c). The spectra recorded with the spectrometer (right after the grating compressor) and FROG are almost the same for the 4 m and 1.6 m fibers and there is no obvious difference in their retrieved spectral phases, as depicted in Figure 5(c) and Figure 7(c), respectively.
It should be noted that although the 4 m fiber produced a broader spectrum after the power amplifier with a spectral width of 38 nm, the corresponding compressed pulse duration was not significantly different from that obtained using the 1.6 m fiber with a spectral width of 19 nm. This is because the long-wavelength components and incoherent ASE present in the 4 m fiber output do not contribute effectively to pulse compression. Moreover, for the 4 m fiber, the spectra measured before and after the gratings showed a noticeable difference. Specifically, the long-wavelength components of the spectrum, or the ASE, appear attenuated after the grating compressor compared to the spectrum before the compressor. This decrease may be attributed to the random polarization of the ASE and the polarization-dependent nature of our transmission gratings (i.e., s-polarized). Essentially, when the gain medium is pumped to achieve population inversion, spontaneous emission occurs in different directions with varying polarizations. Because our gratings preferentially transmit s-polarized light, some wavelengths are diffracted efficiently for one polarization, while others are poorly diffracted. As a result, the ASE appears to be attenuated after the grating compressor.
5 Summary and conclusions
@@
In summary, we have improved our previous CPA system by redesigning the amplification stage and incorporating a CFBG for tunable dispersion control in the compression stage. With the controllable CFBG, the pulse was compressed with reasonable spectral quality, as the spectrogram showed independence from the pumping power, and with an improved temporal profile, as it is free from side artifacts and strong nonlinear effects. We have also highlighted the characteristics of pulses generated using a long fiber with a small core (4 m fiber) and a short fiber with a large core (1.6 m fiber). For both cases, we achieved watt-level average power and transform-limited duration of 210 fs. We found that the 4 m fiber produced a broader amplified spectrum, which was largely associated with long-wavelength ASE that did not contribute to effective pulse compression. In contrast, the 1.6 m fiber provided a higher signal-to-noise ratio, better suppression of ASE, and cleaner amplified spectrum. Although, in general, the 1.6 m fiber delivers improved pulse quality, we think that these comparisons between the 4 m and 1.6 m gain fibers add caution when designing fiber-based amplifier laser system. We also believe this is the first all-Tm:Tb:ZBLAN amplifier system that directly generates watt-level, short-wavelength ultrafast pulses. The construction of the all-Tm:Tb:ZBLAN amplifier system offers a simpler design while still delivering high average power and pulse duration near 200 fs, making it a suitable light source for three-photon lifetime imaging microscopy experiments with red fluorescent proteins [28, 29, 30].
@@
National Institute for Physiological Sciences (24NIPS225, 25NIPS233); National Institute of Natural Sciences (OML0123008,OML012410)